Shift Summary, May 30 DAY shift

Vivian O'Dell (odell)
Sun, 30 May 99 15:40:04 -0500

Shift Summary for Day shift, 05/30/99
On Shift: Albert Erwin
Vivian O'Dell
A variety of other committed collaborators were working throughout this
shift.

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

No beam this shift due to TeV studies.

Activities:

7:00 Show up for shift. Informed "no beam 'till Wednesday"

9:00 Rick T. drops by and works on replacing bad ADC's in crate
14 (they have dead bits). Eventually it works. We start DA to
make sure everything is hunky-dory. Rick notices some peds have
shifted and points it out to Valeri. (RC9/RC10 -- see log book).

11:00 Start Cosmic run, because we can.

11:30 notice DC1Y tripped. Called Hogan, who said he'd come in
later and work on it. Also noticed one of the kumquat relay racks
voltages is low. Call Jason. He says it's no big deal.

13:00 Hogan arrives, brings up DC1Y and says everything is fine.

13:30 In consultation with MCR we dropped interlocks (this time
intentionally). Feel free to stroll the hall for the next couple
of days.

13:30 Val calls & we tell him about bad cap-ids which Hogan has
noticed. Try resetting w/dbc several times, it doesn't help. Page
Julie, who pointed out she had already noticed this last night and
that she suspects timing. She's informed PSh about it.

14:14 RJT adjusts HV on MU2/MU3. He wants a long muon run to calibrate
the muon detectors.

15:30 0.5 of new shift arrives.

15:30 PSh arrives to sort out the capid problem, or find someone
else to blame for it.


Rick T (who remained upstairs), Al replaced the terminator on
the PMT base for MU3 channel 32X which fixed that channel.

7:30 took loooooooooooong laser scan 13215.

7:30 Sydney finds a problem in the onsplit s/w and fixes it.
Another problem he couldn't fixed and sent email to JAG.

10:00 Al realized he forgot his sweater downstairs. He remained
cold the rest of the shift.

12:00 Al made more copies of the gas check list, and checks the
gas. He notes that the alcohol level can no longer be checked because
of the backpressure on the bubbler.

14:00 Arun tests new damp_summary. Mostly works.

14:20 Many CsI channels have shifted quite a bit, according to the
muon run from last night (see list in logbook). Ed installed new
CsI Q/E constants based on muon run 13208.

15:15 Albert and Valeri made an access to adjust HV on V/V'. Albert
retrieves his sweater. Interlocks accidently dropped. Call MCR. They
will send someone out for a search & secure but don't know when.

15:30 new shift takes over.

=========================================================================
Runs ended this shift:
Run 131215 - loooong laser scan

Tapes brought to FCC this shift: 0